A principal of $2500 is invested; part at 8% per annum and the rest at 12% per annum. In one year, the two investments earned equal amounts of interest. How much was invested at each rate?
Answer by checkley71(8403)

.12*x=.08(2500-x) or .12x=200-.08x or .2x=200 or x=200/.2 or x=1000 thus
$1,000*.12=$120 & .08*(2500-1000) or .08*1500=120 thus the interest for both investments are $120.